OpenStudy (anonymous):

A cardboard box of 32 in^3 volume with a square base and open top is to be constructed. What is the length of base that will minimize the surface area?

OpenStudy (dumbcow):

Volume = Area_Base * height 32 = x^2 * h --> h = 32/x^2 Surface area = perimeter_Base*h + Area_Base SA = 4x(32/x^2) + x^2 SA = 128/x + x^2 to minimize, set derivative equal to zero SA' = -128/x^2 + 2x = 0 2x = 128/x^2 2x^3 = 128 x^3 = 64 x = 4 Surface area minimized when length of base is 4 in.